# Attenuation Bias

### v2

@@ 8 Guidance, Limitations, Extensions
The derivations and analyses above illustrate the unique challenges encountered when estimating regression models involving a latent variable as a predictor. We demonstrate that using standard regression models with a latent predictor can yield highly attenuated estimates, whereas MOC, which is likely the most prominent method for correction in political science, typically exacerbates these biases. Standard IV corrections can result in estimates that are even more biased in the opposite direction. Our corrected split-sample IV delivers consistent slope estimates, reducing bias (often dramatically) and performing similarly to more cumbersome approaches such as full joint estimation of the latent traits and regression models. In contrast to other approaches, our method can also be used with any estimation strategy for the latent trait, including those without explicit models or likelihoods.
@@
Applied researchers using latent predictors should check standard estimates against our corrected estimates. In some cases, we may expect attenuation bias to be negligible. For example, estimates of legislator ideology based on hundreds of roll call votes tend to be estimated extremely precisely, at least conditional on the assumed measurement models [32]. NOMINATE legislator ideology scores estimated separately with the first and second sessions of the 117th U.S. Senate are correlated at 0.99, implying that naive OLS regressions using NOMINATE scores would be attenuated by less than one percent. Similarly, the example from Section 7.1 above of democracy predicting economic growth shows less than a 1% change in the estimated coefficient after using our corrected estimator.
@@
The estimated coefficients on other latent predictors may show non-trivial, albeit not huge, changes when applying our corrected methods. For example, [5] finds that correlations between various issue scales when estimated from the two halves of ANES items range from 0.61 to 0.84. This implies that naive OLS slope estimates in these situations may need to be adjusted upwards in magnitude by between 9 and 28 percent. As shown in Section 7.2 above, correcting estimates of the relationship between Congressional speech sentiment and subsequent incumbent vote share produces smaller, but still substantively meaningful changes of around 14% in magnitude.
@@
But in other applications for latent predictors, attenuation bias may be much more severe. It is quite common for latent traits to be estimated based on a very small number of indicators. The political knowledge example presented above demonstrates that correcting naive regression estimates using our proposed method yields a roughly 50% increase in the estimated slope coefficient. Furthermore, our simulations in Section 6 demonstrate that when using 10-20 policy questions to estimate respondents’ ideology, our corrected estimators can result in more than 10-25% increases in the magnitude of estimates compared to naive estimation, and when using 6 indicators, correcting estimates can increase the magnitude of estimates by over one third.
@@
There are several avenues for future work. For example, constructing optimal partitions of indicators for our corrected IV approach could minimize the variability of estimates; in settings where external instruments are unavailable, instruments can sometimes be constructed from heteroskedasticity [31]; these ideas complement our split-indicator strategy. We also plan to explore whether the latent variable case, with its multiple indicators that can be used to construct various sets of IVs, may allow for useful over-identification tests, perhaps of the assumptions necessary for the consistent estimation of the slope coefficient of interest. Finally, we hope to expand our focus on linear models to encompass other regression frameworks (e.g., GLMs) as well as hierarchical models.
@@
$\frac{\beta}{\sqrt{1+\sigma^{2}_{U_{1}}}}\left[(1+\sigma^{2}_{U_{1}})(1+\sigma^{2}_{U_{2}})\right]^{\frac{1}{4}}$ $=\beta\frac{(1+\sigma^{2}_{U_{2}})^{\frac{1}{4}}}{(1+\sigma^{2}_{U_{1}})^{\frac{1}{4}}}.$ (9)
@@
Figure A.I.1: This figure depicts the value of the factor $\frac{(1+\sigma^{2}_{U_{2}})^{\frac{1}{4}}}{(1+\sigma^{2}_{U_{1}})^{\frac{1}{4}}}$ from Equation 9 as a function of $\sigma^{2}_{U_{1}}$ and $\sigma^{2}_{U_{2}}$. The left pane shows the values of this factor when each of these variances ranges between 0 and 1, while the right pane shows values as the variances range between 0 and 10.
But how large does the difference between $\sigma^{2}_{U_{1}}$ and $\sigma^{2}_{U_{1}}$ have to be in order to produce a meaningfully large skew from using this correlation-based correction? In order to assess this Figure A.I.1 plots the factor multiplying $\beta$ in 9 as a function of both measurement error variances. The left pane shows the factor multiplying $\beta$ on the right side of 9 as the measurement error variances for the two estimates of the latent trait, each of which ranges between 0 and 1. For most applications, we would expect ex ante that these parameters would fall into this range, i.e., that the measurement error variance for each of the two sets of estimates would not be larger than the sample variance of the true latent trait itself. Within this range, we observe that the factor is small but not always negligible, reaching values below 0.85 (a 15 percent downward skew in the slope estimate) and above 1.15 (a 15 percent upward skew in the slope estimate) for some instances. But note that these values are ones in which one of the measures has measurement error variance very close to 0, and the other has measurement error variance very close to 1. If we were to partition a set of indicators evenly into two sets, it would seem unlikely that we would be in such a situation. But skewing the slope estimate by smaller amounts, e.g., by 5 or 10 percent in either direction, may still be problematic.
@@
$\nicefrac{{\left[\frac{(1+\sigma^{2}_{U_{2}})^{\frac{1}{4}}}{(1+\sigma^{2}_{U_{1}})^{\frac{1}{4}}}+\frac{(1+\sigma^{2}_{U_{1}})^{\frac{1}{4}}}{(1+\sigma^{2}_{U_{2}})^{\frac{1}{4}}}\right]}}{{2}}.$ (10)
@@
Also note that the expression in Equation 10 holds for a particular partition of the indicators, where the estimates are averaged using one or the other as the predictor. Our proposed corrected estimator $\hat{\beta}^{*}$ takes the median over all such within-partitions averages of the corrected slope estimate, which presumably makes it even less likely that violations of this identical variance assumption would result in large differences in practice.
@@
Figure A.I.2: This plot displays the value of the factor $\nicefrac{{\left[\frac{(1+\sigma^{2}_{U_{2}})^{\frac{1}{4}}}{(1+\sigma^{2}_{U_{1}})^{\frac{1}{4}}}+\frac{(1+\sigma^{2}_{U_{1}})^{\frac{1}{4}}}{(1+\sigma^{2}_{U_{2}})^{\frac{1}{4}}}\right]}}{{2}}$ from Equation 10 as a function of $\sigma^{2}_{U_{1}}$ and $\sigma^{2}_{U_{2}}$. The left pane shows the values of this factor when each of these variances ranges between 0 and 1, while the right pane shows values as the variances range between 0 and 10.
@@
This section illustrates how the standard IV correction for measurement error yields biased upwards in magnitude estimates in the case of latent predictors. It also demonstrates how in the split indicators setup the correlation between $\hat{X}_{1}$ and $\hat{X}_{2}$ can be used to adjust the IV estimator to provide consistent estimates. Finally, it is shown that the corrected IV estimate is equivalent to the corrected OLS estimate in the bivarate case.
One common method for correcting for measurement error in a predictor is the use of instrumental variables (IV). Although its use today is often motivated by causal identification, IV can also, under certain assumptions, produce consistent estimates of slope coefficients when a predictor is measured with error. [^19] Although we focus on measurement error here, note that in a causal framework, IV can be used to address measurement error as well as simultaneity and omitted variable bias under suitable assumptions (see [37]).
@@
Following [4], the two-stage least squares (2SLS) IV estimator of the slope if we were able to observe $\tilde{X}$ could be written as
$\hat{\beta}_{\textrm{IV},\tilde{X}}=\frac{\hat{\beta}_{Y\sim Z}}{\hat{\beta}_{\tilde{X}\sim Z}}=\frac{\nicefrac{{\textrm{Cov}(Y,Z)}}{{\textrm{Var}(Z)}}}{\nicefrac{{\textrm{Cov}(\tilde{X},Z)}}{{\textrm{Var}(Z)}}}=\frac{\textrm{Cov}(X\beta+\varepsilon,Z)}{\textrm{Cov}(X+U,Z)}=\frac{\beta\textrm{Cov}(X,Z)}{\textrm{Cov}(X,Z)}$ (11)
@@
$\operatornamewithlimits{plim}\hat{\beta}_{\textrm{IV},\tilde{X}}=\beta_{X}$ (12)
@@
Because these are both measures of the same latent trait, they should be highly correlated with each other (which we can also check empirically) and with the true latent trait XX. [^20] In the latent predictor case, different indicators indicators of the same latent trait, and especially scales constructed based on multiple indicators, are typically highly correlated with each other[5, see e.g.], suggesting that the weak instruments problem should rarely be severe in practice. We may also assume that the instrument is independent of the measurement error, i.e., that $\hat{X}_{2}\perp\!\!\!\perp U_{1}\equiv(X-\hat{X}_{1})$.
In contrast to the IV setup with an observable predictor, our estimates $\hat{X}_{1}$ and $\hat{X}_{2}$ are not on the same scale as $X$ itself. If we had $\tilde{X}_{1}=X+U_{1}$ to use as our predictor, then the IV results in Equations 11 and 12 would hold. In our case, however, we use as our predictor $\hat{X}_{1}=\frac{\tilde{X}_{1}}{\sqrt{1+\sigma^{2}_{U}}}$, i.e. a rescaled version of $\tilde{X}$ since it is standardized by dividing by $\sigma_{\tilde{X}}=\sqrt{1+\sigma_{U}^{2}}$. [^21] Note that this is a slight abuse of notation since here $U$ is the measurement error in the latent trait estimates based on the first half of the indicators. In other words, what is being referred to as $\tilde{X}$ (and also $\hat{X}$) has changed to now refer to the latent trait estimates based on half of the indicators. This implies that our IV estimate using $\hat{X}$ will have the following relation with $\hat{\beta}_{\textrm{IV},\tilde{X}}$
$\hat{\beta}_{\textrm{IV},\hat{X}_{1}\mid\hat{X}_{2}}=\hat{\beta}_{\textrm{IV},\tilde{X}_{1}\mid\tilde{X}_{2}}\sqrt{1+\sigma_{U}^{2}}$ (13)
@@
$\operatornamewithlimits{plim}\hat{\beta}_{\textrm{IV},\hat{X}_{1}\mid\hat{X}_{2}}=\beta\sqrt{1+\sigma_{U}^{2}},$ (14)
@@
$\hat{\beta}^{*}_{\textrm{IV},\hat{X}_{1}\mid\hat{X}_{2}}=\frac{\hat{\beta}_{\textrm{IV},\hat{X}_{1}\mid\hat{X}_{2}}}{\sqrt{1+\sigma_{U}^{2}}}=\hat{\beta}_{\textrm{IV},\hat{X}_{1}\mid\hat{X}_{2}}\sqrt{\lambda}.$ (15)
@@
$\operatornamewithlimits{plim}\hat{\beta}^{*}_{\textrm{IV},\hat{X}_{1}\mid\hat{X}_{2}}=\beta$ (16)
@@
$\hat{\beta}^{*}_{\textrm{IV},\hat{X}_{1}\mid\hat{X}_{2}}=\hat{\beta}_{\textrm{IV},\hat{X}_{1}\mid\hat{X}_{2}}\;\sqrt{\text{Cor}(\hat{X}_{1},\hat{X}_{2})}.$ (17)
Another way of understanding the corrected IV estimator is to note that if we transformed our measures $\hat{X}_{1}$ and $\hat{X}_{2}$ from having variance 1 to having variance $1+\sigma^{2}_{U}$, which would effectively mean we are in the standard measurement error setup, then the standard IV estimator would provide consistent estimates. [^22] We actually would not need to apply this transformation to whichever one is used as the instrument, since any affine transformation of the instrument will cancel out in the expression for the IV estimator. A convenient way to do this is to estimate $\sigma^{2}_{U}$ with $\nicefrac{{1}}{{\text{Cor}(\hat{X}_{1},\hat{X}_{2})}}-1]$, then note that since we want the variance of our predictor to be $1+\sigma^{2}_{U}$ we can multiply it by $\nicefrac{{1}}{{\sqrt{\text{Cor}(\hat{X}_{1},\hat{X}_{2})}}}$ and use this transformed predictor in a standard IV analysis. [^23] Since the ratio of the uncorrected OLS and uncorrected IV estimators provides information about $\lambda$ (see [37]), yet another equivalent estimator can be constructed based on the square root of the product of the baseline OLS and uncorrected IV estimators using the partitioned $\hat{X}_{1}$ and $\hat{X}_{2}$. Since this approach is less intuitive, we do not discuss it here. Obviously, multiplying the predictor by $\nicefrac{{1}}{{\sqrt{\text{Cor}(\hat{X}_{1},\hat{X}_{2})}}}$ has the result of dividing the estimated slope coefficient by the same factor, which is equivalent to multiplying it by $\sqrt{\text{Cor}(\hat{X}_{1},\hat{X}_{2})}$ as in Equation 17 above.
@@
Finally, the framework here assumes classical (mean-zero, independent) measurement error. When non-classical error is a concern, identification can still sometimes be recovered under additional conditions; see, for example, [23]. In settings where external instruments are unavailable, instruments can sometimes be constructed from heteroskedasticity [31]; these ideas complement our split-indicator strategy.
@@
$\hat{\beta}_{\text{IV},\hat{X}_{1}\mid\hat{X}_{2}}$ $=\frac{\hat{\beta}_{\text{OLS},Y\sim\hat{X}_{2}}}{\hat{\beta}_{\text{OLS},\hat{X}_{1}\sim\hat{X}_{2}}}$ (18)
$=\frac{\hat{\beta}_{\text{OLS},Y\sim\hat{X}_{2}}}{\text{Cor}(\hat{X}_{1},\hat{X}_{2})}$ (19)
@@
$\hat{\beta}^{*}_{\text{IV},\hat{X}_{1}\mid\hat{X}_{2}}$ $=\hat{\beta}_{\text{IV},\hat{X}_{1}\mid\hat{X}_{2}}\;\sqrt{\text{Cor}(\hat{X}_{1},\hat{X}_{2})}.$ (20)
@@
We then set the true values for 1,000 respondents’ ideal points to those point estimates for $X_{i}$ (after imposing the identifying restriction that these ideal points have mean 0, variance 1, and that higher values correspond to more conservative positions) and set the true values of the difficulty ($\eta$) and discrimination parameters ($\alpha$) to those estimated from these indicators in our CCES sample. We also set the true values for the intercept $\beta_{0}$, the slope $\beta_{x}$, and the error variance $\textrm{Var}(\epsilon_{i})$ to roughly correspond to the values estimated in the linear regression predicting presidential vote with the estimated ideal points using the baseline OLS approach. [^26] The estimated slope using our corrected estimates is slightly under 0.4, but for the purposes of the simulations, we use the value of 0.4 since a round number is easier to use as a baseline for comparison in figures and elsewhere. Note that this value for the slope is likely to be attenuated, i.e., smaller in magnitude than the true slope coefficient, since it was estimated from data that itself includes measurement error in the predictor, but we leave it as is for the true value in our simulations.
@@
Finally, the bottom pane of Figure A.III.1 shows split-sample IV estimates corrected using the correlation between $\hat{X}_{1}$ and $\hat{X}_{2}$ as described in Equation 17. These estimates are quite similar to those using the true $X$, but with a very slightly higher bias and RMSE. This suggests the feasibility of estimating $\beta_{X}$ with the corrected IV approach in the more realistic case where we cannot use the true $X$ to estimate the measurement error variance of the split-half estimates.
@@
A.III.1.4 Comparison with Bayesian Strategies
@@
The method of composition approach draws from the posterior distribution of the latent trait estimates [47, 45]. We first fit the outcome OLS model to each posterior draw of the latent traits. We then draw random samples from the resulting likelihood distribution, using draws from a Multivariate Gaussian centered at the point estimate and with a heteroskedasticity-consistent variance-covariance matrix. See [47] for details (especially the right panel of p. 215).
Multiple imputation, described in [41] and refined in [9] for the joint outcome and measurement error modeling case, involves generating multiple plausible values for the latent traits using mean and variances from an initial MCMC step as priors—analyzing each imputed dataset separately before combining results.

### v3

@@ 6 Guidance, Limitations, Extensions
The derivations and analyses above illustrate the unique challenges encountered when estimating regression models involving a latent variable as a predictor. We demonstrate that using standard regression models with a latent predictor can yield highly attenuated estimates, whereas MOC, which is likely the most prominent method for correction in political science, typically exacerbates these biases. Standard IV corrections can result in estimates that are even more biased in the opposite direction. Our correlation-corrected estimator reduces bias in slope estimates, often dramatically, and performs similarly to more cumbersome approaches such as full joint estimation of the latent traits and regression models. In contrast to other approaches, our method can also be used with any estimation strategy for the latent trait, including those without explicit models or likelihoods.
@@
Applied researchers using latent predictors should check standard estimates against our corrected estimates. In some cases, we may expect attenuation bias to be negligible. For example, estimates of legislator ideology based on hundreds of roll call votes tend to be estimated extremely precisely, at least conditional on the assumed measurement models [undefab]. NOMINATE legislator ideology scores estimated separately with the first and second sessions of the 117th U.S. Senate are correlated at 0.99, implying that standard OLS regressions using NOMINATE scores would be attenuated by less than one percent. Appendix A.IV.1.6 presents an application of democracy predicting economic growth that shows less than a 1% change in the estimated coefficient after using our corrected estimator.
@@
The estimated coefficients on other latent predictors may show non-trivial, albeit not huge, changes when applying our corrected methods. For example, [undefd] finds that correlations between various issue scales when estimated from the two halves of ANES items range from 0.61 to 0.84. This implies that standard OLS slope estimates in these situations may need to be adjusted upwards in magnitude by between 9 and 28 percent. As shown in Appendix IV, correcting estimates of the relationship between Congressional speech sentiment and subsequent incumbent vote share produces smaller, but still substantively meaningful changes of around 14% in magnitude.
@@
But in other applications for latent predictors, attenuation bias may be much more severe. It is quite common for latent traits to be estimated based on a very small number of indicators. The political knowledge example presented above demonstrates that correcting standard regression estimates using our proposed method yields a roughly 50% increase in the estimated slope coefficient.
Furthermore, our simulations show how correction can matter in practice. With 10–20 policy questions for ideology, the corrected estimates are often 10–25% larger than standard estimates. When using 6 indicators, correcting estimates can increase their magnitude by over one-third.
@@
There are several avenues for future work. One extension is to choose indicator splits that reduce the variance of estimates. Also, when split samples cannot be constructed, researchers can sometimes build instruments from heteroskedasticity. Both ideas fit naturally with our split-indicator approach. We also plan to explore whether the latent variable case, with its multiple indicators that can be used to construct various sets of IVs, may allow for useful over-identification tests, perhaps of the assumptions necessary for the consistent estimation of the slope coefficient of interest. Finally, we hope to expand our focus on linear models to encompass other regression frameworks such as GLMs. $\blacksquare$
@@
$\text{Cor}(\hat{X}_{1},\hat{X}_{2})$ $=\text{Cor}(\tilde{X}_{1},\tilde{X}_{2})=\frac{\text{Cov}(\tilde{X}_{1},\tilde{X}_{2})}{\sqrt{\sigma^{2}_{\tilde{X}_{1}}\sigma^{2}_{\tilde{X}_{2}}}}$ (A.I.1)
@@
$\frac{\beta}{\sqrt{1+\sigma^{2}_{U_{1}}}}\left[(1+\sigma^{2}_{U_{1}})(1+\sigma^{2}_{U_{2}})\right]^{\frac{1}{4}}$ $=\beta\frac{(1+\sigma^{2}_{U_{2}})^{\frac{1}{4}}}{(1+\sigma^{2}_{U_{1}})^{\frac{1}{4}}}.$ (A.I.2)
@@
Figure A.I.1: This figure depicts the value of the factor $\frac{(1+\sigma^{2}_{U_{2}})^{\frac{1}{4}}}{(1+\sigma^{2}_{U_{1}})^{\frac{1}{4}}}$ from Equation A.I.2 as a function of $\sigma^{2}_{U_{1}}$ and $\sigma^{2}_{U_{2}}$. The left pane shows the values of this factor when each of these variances ranges between 0 and 1, while the right pane shows values as the variances range between 0 and 10.
But how large does the difference between $\sigma^{2}_{U_{1}}$ and $\sigma^{2}_{U_{1}}$ have to be in order to produce a meaningfully large skew from using this correlation-based correction? In order to assess this Figure A.I.1 plots the factor multiplying $\beta$ in Equation A.I.2 as a function of both measurement error variances. The left pane shows the factor multiplying $\beta$ on the right side of Equation A.I.2 as the measurement error variances for the two estimates of the latent trait, each of which ranges between 0 and 1. For most applications, we would expect ex ante that these parameters would fall into this range, i.e., that the measurement error variance for each of the two sets of estimates would not be larger than the sample variance of the true latent trait itself. Within this range, we observe that the factor is small but not always negligible, reaching values below 0.85 (a 15 percent downward skew in the slope estimate) and above 1.15 (a 15 percent upward skew in the slope estimate) for some instances. But note that these values are ones in which one of the measures has measurement error variance very close to 0, and the other has measurement error variance very close to 1. If we were to partition a set of indicators evenly into two sets, it would seem unlikely that we would be in such a situation. But skewing the slope estimate by smaller amounts, e.g., by 5 or 10 percent in either direction, may still be problematic.
@@
$\nicefrac{{\left[\frac{(1+\sigma^{2}_{U_{2}})^{\frac{1}{4}}}{(1+\sigma^{2}_{U_{1}})^{\frac{1}{4}}}+\frac{(1+\sigma^{2}_{U_{1}})^{\frac{1}{4}}}{(1+\sigma^{2}_{U_{2}})^{\frac{1}{4}}}\right]}}{{2}}.$ (A.I.3)
@@
Also note that the expression in Equation A.I.3 holds for a particular partition of the indicators, where the estimates are averaged using one or the other as the predictor. Our proposed corrected estimator $\hat{\beta}^{*}$ takes the median over all such within-partitions averages of the corrected slope estimate, which presumably makes it even less likely that violations of this identical variance assumption would result in large differences in practice.
@@
Figure A.I.2: This plot displays the value of the factor $\nicefrac{{\left[\frac{(1+\sigma^{2}_{U_{2}})^{\frac{1}{4}}}{(1+\sigma^{2}_{U_{1}})^{\frac{1}{4}}}+\frac{(1+\sigma^{2}_{U_{1}})^{\frac{1}{4}}}{(1+\sigma^{2}_{U_{2}})^{\frac{1}{4}}}\right]}}{{2}}$ from Equation A.I.3 as a function of $\sigma^{2}_{U_{1}}$ and $\sigma^{2}_{U_{2}}$. The left pane shows the values of this factor when each of these variances ranges between 0 and 1, while the right pane shows values as the variances range between 0 and 10.
@@
This section illustrates how the standard IV correction for measurement error yields biased upwards in magnitude estimates in the case of latent predictors. It also demonstrates how in the split indicators setup the correlation between $\hat{X}_{1}$ and $\hat{X}_{2}$ can be used to adjust the IV estimator to provide consistent estimates. Finally, it is shown that the correlation-corrected IV estimate is equivalent to the correlation-corrected OLS estimate in the bivarate case.
One common method for correcting for measurement error in a predictor is the use of instrumental variables (IV). Although its use today is often motivated by causal identification, IV can also, under certain assumptions, produce consistent estimates of slope coefficients when a predictor is measured with error. [^15] Although we focus on measurement error here, note that in a causal framework, IV can be used to address measurement error as well as simultaneity and omitted variable bias under suitable assumptions (see [undefae]).
@@
Following [undefc], the two-stage least squares (2SLS) IV estimator of the slope if we were able to observe $\tilde{X}$ could be written as
$\hat{\beta}_{\textrm{IV},\tilde{X}}=\frac{\hat{\beta}_{Y\sim Z}}{\hat{\beta}_{\tilde{X}\sim Z}}=\frac{\nicefrac{{\textrm{Cov}(Y,Z)}}{{\textrm{V}(Z)}}}{\nicefrac{{\textrm{Cov}(\tilde{X},Z)}}{{\textrm{V}(Z)}}}=\frac{\textrm{Cov}(X\beta+\varepsilon,Z)}{\textrm{Cov}(X+U,Z)}=\frac{\beta\textrm{Cov}(X,Z)}{\textrm{Cov}(X,Z)}.$ (A.II.1)
@@
$\operatornamewithlimits{plim}\hat{\beta}_{\textrm{IV},\tilde{X}}=\beta_{X}.$ (A.II.2)
@@
Because these are both measures of the same latent trait, they should be highly correlated with each other, which we can also check empirically, and also highly correlated with the true latent trait XX. [^16] In the latent predictor case, different indicators indicators of the same latent trait, and especially scales constructed based on multiple indicators, are typically highly correlated with each other [undefd, see e.g.], suggesting that the weak instruments problem should rarely be severe in practice. We may also assume that the instrument is independent of the measurement error, i.e., that $\hat{X}_{2}\perp\!\!\!\perp U_{1}\equiv(X-\hat{X}_{1})$. [^17] This is analogous to the conditional independence assumption used in many latent traits models, i.e. that conditional on $X$ the indicators are independent of each other.
In contrast to the IV setup with an observable predictor, our estimates $\hat{X}_{1}$ and $\hat{X}_{2}$ are not on the same scale as $X$ itself. If we had $\tilde{X}_{1}=X+U_{1}$ to use as our predictor, then the IV results in Equations A.II.1 and A.II.2 would hold. In our case, however, we use as our predictor $\hat{X}_{1}=\frac{\tilde{X}_{1}}{\sqrt{1+\sigma^{2}_{U}}}$, i.e. a rescaled version of $\tilde{X}$ since it is standardized by dividing by $\sigma_{\tilde{X}}=\sqrt{1+\sigma_{U}^{2}}$. [^18] Note that this is a slight abuse of notation since here $U$ is the measurement error in the latent trait estimates based on the first half of the indicators. In other words, what is being referred to as $\tilde{X}$ (and also $\hat{X}$) has changed to now refer to the latent trait estimates based on half of the indicators. This implies that our IV estimate using $\hat{X}$ will have the following relation with $\hat{\beta}_{\textrm{IV},\tilde{X}}$
$\hat{\beta}_{\textrm{IV},\hat{X}_{1}\mid\hat{X}_{2}}=\hat{\beta}_{\textrm{IV},\tilde{X}_{1}\mid\tilde{X}_{2}}\sqrt{1+\sigma_{U}^{2}}$ (A.II.3)
@@
$\operatornamewithlimits{plim}\hat{\beta}_{\textrm{IV},\hat{X}_{1}\mid\hat{X}_{2}}=\beta\sqrt{1+\sigma_{U}^{2}},$ (A.II.4)
@@
$\hat{\beta}^{*}_{\textrm{IV},\hat{X}_{1}\mid\hat{X}_{2}}=\frac{\hat{\beta}_{\textrm{IV},\hat{X}_{1}\mid\hat{X}_{2}}}{\sqrt{1+\sigma_{U}^{2}}}=\hat{\beta}_{\textrm{IV},\hat{X}_{1}\mid\hat{X}_{2}}\sqrt{\lambda}.$ (A.II.5)
@@
$\operatornamewithlimits{plim}\hat{\beta}^{*}_{\textrm{IV},\hat{X}_{1}\mid\hat{X}_{2}}=\beta,$ (A.II.6)
@@
$\hat{\beta}^{*}_{\textrm{IV},\hat{X}_{1}\mid\hat{X}_{2}}=\hat{\beta}_{\textrm{IV},\hat{X}_{1}\mid\hat{X}_{2}}\;\sqrt{\text{Cor}(\hat{X}_{1},\hat{X}_{2})}.$ (A.II.7)
Another way of understanding the corrected IV estimator is to note that if we transformed our measures $\hat{X}_{1}$ and $\hat{X}_{2}$ from having variance 1 to having variance $1+\sigma^{2}_{U}$, which would effectively mean we are in the standard measurement error setup, then the standard IV estimator would provide consistent estimates. [^19] We actually would not need to apply this transformation to whichever one is used as the instrument, since any affine transformation of the instrument will cancel out in the expression for the IV estimator. A convenient way to do this is to estimate $\sigma^{2}_{U}$ with $\nicefrac{{1}}{{\text{Cor}(\hat{X}_{1},\hat{X}_{2})}}-1]$, then note that since we want the variance of our predictor to be $1+\sigma^{2}_{U}$ we can multiply it by $\nicefrac{{1}}{{\sqrt{\text{Cor}(\hat{X}_{1},\hat{X}_{2})}}}$ and use this transformed predictor in a standard IV analysis. [^20] Since the ratio of the uncorrected OLS and uncorrected IV estimators provides information about $\lambda$ (see [undefae]), yet another equivalent estimator can be constructed based on the square root of the product of the baseline OLS and uncorrected IV estimators using the partitioned $\hat{X}_{1}$ and $\hat{X}_{2}$. Since this approach is less intuitive, we do not discuss it here. Obviously, multiplying the predictor by $\nicefrac{{1}}{{\sqrt{\text{Cor}(\hat{X}_{1},\hat{X}_{2})}}}$ has the result of dividing the estimated slope coefficient by the same factor, which is equivalent to multiplying it by $\sqrt{\text{Cor}(\hat{X}_{1},\hat{X}_{2})}$ as in Equation A.II.7 above.
@@
Finally, the framework here assumes classical (mean-zero, independent) measurement error. When non-classical error is a concern, identification can still sometimes be recovered under additional conditions; see, for example, [undefu]. In settings where external instruments are unavailable, instruments can sometimes be constructed from heteroskedasticity [undefaa]; these ideas complement our split-indicator strategy.
@@
$\hat{\beta}_{\text{IV},\hat{X}_{1}\mid\hat{X}_{2}}$ $=\frac{\hat{\beta}_{\text{OLS},Y\sim\hat{X}_{2}}}{\hat{\beta}_{\text{OLS},\hat{X}_{1}\sim\hat{X}_{2}}}$ (A.II.8)
$=\frac{\hat{\beta}_{\text{OLS},Y\sim\hat{X}_{2}}}{\text{Cor}(\hat{X}_{1},\hat{X}_{2})}$ (A.II.9)
@@
$\hat{\beta}^{*}_{\text{IV},\hat{X}_{1}\mid\hat{X}_{2}}$ $=\hat{\beta}_{\text{IV},\hat{X}_{1}\mid\hat{X}_{2}}\;\sqrt{\text{Cor}(\hat{X}_{1},\hat{X}_{2})}.$ (A.II.10)
@@
We then set the true values for 1,000 respondents’ ideal points to those point estimates for $X_{i}$ (after imposing the identifying restriction that these ideal points have mean 0, variance 1, and that higher values correspond to more conservative positions) and set the true values of the difficulty ($\eta$) and discrimination parameters ($\alpha$) to those estimated from these indicators in our CCES sample. We also set the true values for the intercept $\beta_{0}$, the slope $\beta_{x}$, and the error variance $\textrm{V}(\epsilon_{i})$ to roughly correspond to the values estimated in the linear regression predicting presidential vote with the estimated ideal points using the baseline OLS approach. [^23] The estimated slope using our corrected estimates is slightly under 0.4, but for the purposes of the simulations, we use the value of 0.4 since a round number is easier to use as a baseline for comparison in figures and elsewhere. Note that this value for the slope is likely to be attenuated, i.e., smaller in magnitude than the true slope coefficient, since it was estimated from data that itself includes measurement error in the predictor, but we leave it as is for the true value in our simulations.
@@
Finally, the bottom pane of Figure A.III.1 shows split-sample IV estimates corrected using the correlation between $\hat{X}_{1}$ and $\hat{X}_{2}$ as described in Equation A.II.7. These estimates are quite similar to those using the true $X$, but with a very slightly higher bias and RMSE. This suggests the feasibility of estimating $\beta_{X}$ with the corrected IV approach in the more realistic case where we cannot use the true $X$ to estimate the measurement error variance of the split-half estimates.
@@
A.III.1.4 Comparison with Additional Estimation Strategies
@@
The method of composition approach draws from the posterior distribution of the latent trait estimates [undefao, undefam]. We first fit the outcome OLS model to each posterior draw of the latent traits. We then draw random samples from the resulting likelihood distribution, using draws from a Multivariate Gaussian centered at the point estimate and with a heteroskedasticity-consistent variance-covariance matrix. See [undefao] for details (especially the right panel of p. 215).
Multiple overimputation, described in [undefah] and refined in [undefh] for the joint outcome and measurement error modeling case, involves generating multiple plausible values for the latent traits using mean and variances from an initial MCMC step as priors—analyzing each imputed dataset separately before combining results.